Companies sometimes consider stock splits to bring down the price so that the stock attracts more purchases.

Consider the following case:

Tolbotics Inc. currently has 20,000 shares of common stock outstanding. Its management believes that its current stock price of $110 per share is too high. The company is planning to conduct stock splits in the ratio of 2 for 1 as described in the animation.

If Tolbotics Inc. declares a 2-for-1 stock split, the price of the company’s stock after the split, assuming that the total value of the firm’s stock remains the same after the split, will be   .

Fuzzy Muffin Manufacturing Company is one of Tolbotics’s leading competitors. Fuzzy Muffin Manufacturing Company’s market intelligence research team shares Tolbotics’s plans of announcing a stock split, influencing the distribution policy makers. Consequently, executives at Fuzzy Muffin decide to offer stock dividends to its shareholders.

A stock dividend is another way of keeping the stock price from going too high. Fuzzy Muffin currently has 2,300,000 shares of common stock outstanding.

If the firm pays a 4% stock dividend, how many shares will the firm issue to its existing shareholders?

101,200 shares

78,200 shares

64,400 shares

92,000 shares
Answer #1

Question 1:

Stock split implies division of existing stock in a defined proportion such that the number of outstanding stocks increase. However, market value of the company does not change as a result of this event.

Tolbotics Inc. declares a 2-for-1 stock split. This implies for every 1 share outstanding, there would be 2 stocks outstanding post split. Therefore, post-split, number of shares outstanding = 20000 * 2 = 40000.

Market price pre split * Number of stocks pre split = Market price post split * Number of stocks post split

20000 * $110 = 40000 * Market price post split

Market price post split = $55

Question 2

Number of share post stock dividend = Number of shares pre-stock dividend * (1 + Stock dividend%)

Number of share post stock dividend = 2,300,000 * (1 + 4%) = 2,392,000

Number of additional shares = 2,392,000 - 2,300,000 = 92,000 shares